OpenStudy (anonymous):

Find the 30th term in the sequence 7, 16, 25, 34, . . . 261 277 270 268

OpenStudy (jhannybean):

Take the preceding term and subtract it by the previous term, you will find a common difference. 16 - 7 =? 25-16 = ? 34- 25 =?

OpenStudy (jhannybean):

This is an arithmetic Progression that follows the format \[\large a_{n} = a_{1} +(n-1)d\] \[\large {a_{30} = 30th\ term \\ a_{1} = 7 \\ n =30 \\ d=?}\]

OpenStudy (jhannybean):

Awesome :)

OpenStudy (anonymous):

np